OpenStudy (anonymous):

2x + 4=8 so what does x=

OpenStudy (anonymous):

realy need help

OpenStudy (mathstudent55):

To solve an equation like this one, look at the operations done to x and undo them in the reverse order. If you had a number to put in for x, to calculate the left side you'd first multiply the number by 2 then you'd add 4 to it. You want to undo all that, so start with the last operation, and undo it. How do you undo + 4? You subtract 4 from both sides. Then how do you undo 2 times? You divide both sides by 2.

OpenStudy (mathstudent55):

Start by subtracting 4 from both sides. Can you write what you get?

OpenStudy (anonymous):

umm lets see

OpenStudy (anonymous):

Which two numbers do I subtract 4 from?

OpenStudy (mathstudent55):

This is how you do it: 2x + 4 - 4 = 8 - 4 I subtracted 4 from both sides. You see -4 on both sides. Now what is 2x + 4 - 4? What is 8 - 4?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so that means that x=4

OpenStudy (mathstudent55):

2x +4 - 4 = 2x 8 - 4 = 4 So you get 2x = 4

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

Now you need to do the next step which is to divide both sides by 2, since the x is being multiplied by 2. 2x/2 = 4/2 What is 2x/2? What is 4/2?

OpenStudy (mathstudent55):

2x/2 = x 4/2 = 2 So x = 2 Now check 2x + 4 = 8, plug in 2 for x and calculate: 2(2) + 4 = 8? 4 + 4 = 8? 8 = 8 Checks, so x = 2 is the solution.

OpenStudy (anonymous):

ok